Minutes July2011 to Dec 2011
11/30/2011: Present: Jixie, Min, Kalyan, Pengjia, Toby, Ryan, Melissa, Jie, JP, Karl,Vince, Alex, Aiden
- JP, Karl:
1) Don will do another cool down at UVA from Dec 8 to 16. Toby will be there to help for the whole
period. Chao will go there and help from 8 to 12 and Jie will help from 12 to 16.
2) We will have the analysis meeting every day at the meeting room located in the 2nd floor
of the counting house from next Monday till the end of the commissioning.
3) Karl sign each component with manpower, check the wiki for details
11/23/2011: Present: Jixie, Min, Kalyan, Pengjia, Toby, Jie, JP, Karl
- Karl and JP:
1) Commissioning will take place from 12/12 to 12/22, straight through beam, C12 target. All subsystems
are supposed to be ready by 12/9.
2) Software for BPM will not be ready for this commissioning. Old SCE readout electronics of the BPM will
be installed. They are supposed to work at beam current no lower than 100nA.
3) Chicane, target and local dump will not be tested.
4) About shift: cancel all shift in Jan. and early Feb. Allow user to sign up shift for the real run.
1) In the irradiation event in NIST, 2-3 cups of target material have been irradiated. Don plan another cool
down at UVA at early December.
2) The hall b magnet is already in the machine shop, waiting for machining.
1) Gave a status update of the scaler and Intel CPU: see details [d here].
- Min and Jixie:
1) Gave a update of the "sieve to target" reconstruct. See details here.
2) Min will provide the self consistency check of this reconstruction for single beam energy version and also
combined beam energy version next week.
3) Bremsstrahlung damage study: Jixie got the equivalent damage tables for n,pr,pi and e- from FLUKA. Will
implement them into the Geant4 program. Here is a plot of the damage as a function of momentum. Check here for more details.
1) Learning the G4 program.
- Chao and Kalyan:
1) Started work on magnet shield for 3rd arm PMTs, waiting for rubber collar to fix the inner tubes
2) Hardware for DAQ is almost done. Working on decoding. Will work on online monitor scripts later.
- Ryan, Melissa:
1) An update of the HRS status can be found here.
11/16/2011: Present: Jixie, Melissa, Min, Kalyan, Alex, Ryan, Jie, JP, Karl
Pengjia, Chao and Toby went to NIST helping target irradiation.
1) In the coming commissioning, 1,3 or 5 passes beam will be available for Hall A.
The beam energy is about 580 MeV per pass, we prefer to 3 pass beam.
2) Need to write the TOSP for the commissioning.
1) Working on the HV crate.
2) Bogdan suggested to replace all the TDC to 1877s. Need to evaluate how much work and how much risk
it is to implement this upgrade.
- Jixie, Min :
1) HRSMC v0.99 released. In this version, HallB magnet geometry and field is built in. Full size
scattering chamber is built. The local dump stand is also in position.
2) Talked to Jack about measuring the septa field. Jack still working on the power supply, might be
ready next week. Jixie need to get the TOSP ready as soon as possible.
3) Finished checking the tosca map with the fortran routine for the HallB target magnet. They match each
other very well.
4) Studied the homogeneity of the hallb magnet. See the figure here.
Compare this coil to the SANE coil, see plot here.
5) HallB field data map for snake is ready. Download it here.
6) Working on a new model for sieve to target reconstruction. The one from MUDIFI is not good at all.
Trying to improve it with new fits.
7) We would also like to have optics runs in the commissioning. Those runs will be used to improve the
reconstruction based on MC data. Need to know the target position before hand.
1) Learning to run the G4 program.
- Ryan, Melissa:
1) EDTM for left arm is working now.
2) Run left HRS in buffer mode. Found 40% dead time. Will investigate this issue in the rest of this week.
11/09/2011: Present: Jixie, Melissa, Min, Kalyan, Alex, Vince, Ryan, Jie
Pengjia, Chao and Toby went to NIST helping target irradiation. Alex is still working on the commissioning plan.
1) Prepare to move 3rd arm DAQ behind the green wall
2) Setting up the helicity scaler gates
3) Decided to use the beam line HV crate
4) Making magnetic shielding for PMTs
5) Trying to improve the reconstruction matrix from sieve to target
- Jixie, Min:
1) Posted the optics discussion and TODO list in [Elog https://hallaweb.jlab.org/dvcslog/g2p/136]
2) Put HallB coil geometry into the G4 model, field module also ready to use.
3) Working on compare the TOSCA model with the routine from Peter Bosted
4) Running simulation for GEP to determine the magnetic angle, (Urgent! need to finish by Thursday)
5) The bremsstrahlung result can be found here
1) Learning to run the Geant4 program
- Vince, Ryan,Melissa:
1) Jack finish install the 3rd fast-bus crate for left arm. Ryan updated the crate map and
took cosmic run to check left HRS DAQ
2) Melissa continue her work on identify the source of low efficiency of right shower. She found that there is 1~2 blocks of shower give low signal. She will spend more time to look into this. 3) Vince, Alex, Ryan will meet with Bob in Thursday to plan a DAQ rate test. This test is trying to
learn how much will be the improvement after install the 3rd fast-bus crate. They might test it again once
Pengjia finished installing the HAPPEX crate into the left HRS.
11/02/2011: Present: Jixie, Melissa, Min, Chao, Pengjia, Kalyan, Alex, Vince, Toby, JP, Aidan, Ryan, Jie, Yawei Zhang
1) Got permission to use Hall-B magnet, will be ready around mid Feb.
2) Will have commissioning from 12/9 to 12/24
3) will work out the commissioning plan then determine the shift schedule
4) Need to finalize the GEP angle. 7 +/- 2 degrees are all acceptable.
Jixie will put the geometry into the G4 model and help to finalize that
- Kalyan, Chao:
1) Find out a way to disable any channel of the ADC for the 3rd arm DAQ. After disable some unused channels,
The DAQ rate can go to 9k Hz in the pulse test.
2) Found the pedestal threshold for ADC
- Jixie, Min, Chao:
1) Got some Hall-B coil drawings, also got a fortran routine to calculate the field from Peter Bosted.
Click here to download them.
2) Will add the geometry and field to the Geant4 modle
3) Will have optics meeting tomorrow from 2:30 to 3:30 right after target meeting at A110. All is welcome to join.
4) Finalize the synchrotron radiation at Chicane: when the photons energy reach 50keV, the rate drops below 1Hz. The
photos came form the 1st magnet will not be able to hit the target because they are blocked by the 2nd magnet. Those
photons came from the 2nd magnet will not hit the target neither.
5) Still working on the bremsstrahlung radiation within the scattering chamber
6) Min is running the simulation for optics
- Pengjia, Jie:
1) working on installing the intel CPU with HSR DAQ
2) Jie is test the ADC of HAPPPEX DAQ for Left and Right HRS
1) Checking preshower|shower to make sure it works.
1) Work with target group. Took HallB magnet chamber down and shipped to EEL building. Will cut it and take the
magnet out soon.
2) continue learning the G4
1) check the preshower|shower using Xgt2 data. Found the one in right HRS have low efficiency(as low as 97%). Vince mentioned
this could be due to the low electron momentum (~0.8 GeV). Melissa will check it with high energy data set.
1) Jack is working on adding a third Fastbus crate for left HRS. When he finishes, Vince will work on checking the left HRS DAQ with the new crate.
10/26/2011: Present: Karl, Jixie, Melissa, Min, Chao, Pengjia, Kalyan, Alex, Vince, Toby, JP, Aidan, Ryan
1) There is an update to the G2P schedule. The start date of this experiment could be middle of Feb (not official date yet).
We have to give up the 12.5 degree runnings.
1) Discussed the g2p commissioning before the target is ready. We can do a lot of detector and DAQ test.
2) There will be 2~3 irradiation carried out at NIST. 3 students will go to learn and help. Pengjia and Toby
will go. There 3 person could be rotated among students. Chao volunteer to go.
- Kalyan, Chao:
1) Test the 3rd arm detector with source (Cs137, average momentum is 150keV, maximum about 500keV).
2) 3rd arm stand is assembled in the hall.
3) keep working to the DAQ, need to find a way to subtract the pedestal in VME.
4) Chao report the calibration
- Jixie, Min:
1) Built chicane into the Geant4 simulation, see 3D plots here
2) Show the study about the synchrotron radiation at chicane. Here is a document about the details of synchrotron.
According to our calculation, the radiation power of all beam energy at chicane magnet B1 is about 20~30 micro-watt.
Jp suggested to calculate the rates for high energy photons which will tell if it will affect the physics result.
3) Still working on the bremsstrahlung radiation at target chamber
4) Min reported the optics status.
5) Pengjia told that we can use the slow raster to change the beam position during cross scan for optic runs. This will
have only 5 minutes of change over, which can save lot of time. JP suggested to check with Yves Roblin about that.
1) Put ADC back to happex daq with voltage mode.
- Ryan, Melissa, Vince:
1) Working on the HRS DAQ.
2) Bogdan agreed to fix the bad PMTs on the left HRS Cherenkov detector
3) Bogdan requested 2 plots for lead glass: "pedestal vs channel" and "cosmic peak vs channel". Ryan will provide them.
4) Vince set up a separate DB for the Right HRS CODA, which solves the platform interference problem.
1) Working with target group
2) Learning the G4 program
10/19/2011: Present: Karl, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, Alex, Vince, Toby, JP, Yuxiang Zhao, Yawei Zhang, Yi Zhou
1) Update of the target: 2 places of short to ground found. Will need to open the coil package for deep
diagnostics. 2 more weeks of delay is expected if we plan to fix the short to ground problem.
2) Looking for feasibility of using the magnet of Hall B. There is lot of constrain to use this magnet. A delay of
2 month will be expected if use it.
1)Readiness and safety review will take place next Monday, 1 PM. Everybody are welcomed to attend and
support us. Check here for details.
1) Presented the scaler map
2) Determined the beam incline angles at target center for various settings using Geant4 program HRSMC with
NEW target field. The vertical and horizontal offset of the beam at 1.1 meter away from the target center are also
extracted. Check here for figures.
1) Presented some figures prepared for the readiness review, including beam spot at local dump, 'flame of sheet', etc...
1) Re-did the analysis for the HRS gas Cherenkovs. Three channels in the left arm seem to have a low number of photoelectrons,
check here for details. Need to report to Bogdan and request for replacement.
1) Reported the HAPPEX DAQ status. Check his talk for details.
1) Figured out the reason for CODA interference of left HRS and right HRS. The platform uses an account name to identify
the database. Both the left and right HRS use the same database name: adaq. Therefore killing the right HRS platform will also
disconnect the left HRS run control. The solution is to change the user account for the platform to run the right HRS DAQ.
10/12/2011: Present: Karl, Jixie, Jie, Min, Chao, Pengjia, Kalyan, Vince, Toby, JP, Yuxiang Zhao
1) In order to improve students' skill of doing presentation, we will have 2-3 students each week to give
feature presentation about what they are doing in the part weeks.
2) Latest schedule can be find here
1) Target status update:
A) There is enough space to put the sodium pad into the coil container
B) Found about 2~3 short grounded place in the coil. Not clear if can fix them all. Waiting for suggestion from Oxford
C) Try to get Hall B Helmholtz coil and make it work for the experiment in the meanwhile (2 magnet are prepare in parallel).
2) Running schedule have been officially changed to Dec. 1st
3) 2-channeled BPM has been tested in hall C and works now. 4-channeled BPM is expected to finish design in 2 more weeks.
- Jixie, Min:
1) A preliminary draft of optics plan can be found here, still need to run more simulation to check some items
2) Jixie is working on the simulation for bremsstrahlung photons and other safety review related items
- Vince, Ryan, Melissa:
1) Ryan gave a nice report of the DAQ status, check here
- Pengjia, Jie:
1) optimizing the HAPPEX ADC module
2) The "buffer full" issue in the HAPPEX DAQ is suddenly gone after recalling. The reason is not figured out yet.
1) Finished the scaler
2) Worked with Pengjia in optimizing the TDC which is using an Intel CPU
3) Learning the geant4 program
- Kalyan, Chao:
1) Test E plane PMTs with cosmic rays, check the details here
10/05/2011: Present: Karl, James, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, Aiden, Vince, Toby, JP, Yuxiang Zhao, Yawei Zhang, Yi Zhou, Zhao Zhang
1) Welcome James joining g2p! James will working on the target related projects for this experiment
2) Latest schedule is available here
3) Encourage all users to upload documentation to this location
1) Got permission from Oxford that the target coil can be cut and opened for investigation. The coil
is now in machine shop.
2) Will work on target operator training procedure and put it on the wiki. Will also arrange target
operator training sessions.
- Kalyan, Chao:
1) Finished gain match (HV adjusting such that each PMT will have the same peak location) for dE planes using
cosmic ray, will do this to the E planes.
2) Try to optimize the DAQ rate to reach 10k Hz.
3) Will do the OSP for the 3rd arm.
4) Show the dE planes gain match result here.
- Jixie, Min:
1) Run elastic simulation with C12, CH2, LHe and Al targets for all possible combination configurations of
beam energies, target field and HRS angles. Calculate the elastic rates for w/o sieve. Also calculate rates for
some single sieve holes. Check the details here.
2) The 1st draft of the optics plan is here.
In this draft, no changing time is considered yet. Will improve this draft and update it.
- Vince, Melissa, Ryan, Toby:
1) Problem of CODA script: killcoda on adaql1 will also cause the run on adaql2 to end, need to find the reason
and solve it.
2) After installing an Intel CPU in ROC 1, the right HRS does not get any triggers. Vince will help Ryan and Pengjia
investigate it after the meeting.
3) Ryan reported the S1 bad PMTs to Bogdan. Some more figures were requested to prove these PMTs are really bad.
He will generate figures for that.
4) Melissa will redo the fit to Cherenkov NPHE with new fitting function (Gaussian+Poisson). She has also redone the
fits for the SPE. With the improved analysis the NPHE will be higher than previous fit.
5) Toby is working on scalers. He needs to verify their connections to both HRS and make sure they work including the
1) The 2nd version of field analysis note is ready for review.
1) Connecting HAPPEX crate to Left HRS, (not finish yet)
9/28/2011: Present: Karl, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, Alex, Vince, Toby, JP, Yuxiang Zhao, Yawei Zhang
1) Target coil quench protection part will be opened today or tomorrow for investigation.
- Kalyan, Chao:
1) 3rd arm daq: have 1 ADC and 1 TDC working so far. Working on the scaler right now. Will make a copy
of ADC and TDC for backup.
2) Tested the dE plane ADC signal with cosmic ray. The trigger is the coincidence of 2 dE planes. The
figure shows that PMTs in the left
have narrower signal compared to those in the right. Need to check reasons for that.
- Jixie, Min:
1) Updated the Gean4 program HRSMC to support titled beam simulation. A uniform field model have been
coded to correct the incident angle of the incoming electron. The effective scattering angle will be
calculated and the elastic scattering energy then will be determined based on the effective scattering
angle. Here is a picture of current version.
2) Simulated the elastic events with C12, CH2 and liquid Helium target with sieve slit in position.
50 mil, 100 mil and 200 mil thick of C12 have been simulated. Check details here.
The result shows that for 1.2 GeV beam, elastic peak of C12 will be overlapped with that of helium.
3) Considering the fact that the uncertainty of deltaP between H peak and C peak is dominated by H peak,
(in other words, the overlap of C peak and He peak will not be the major source of uncertainty), we
will not warn up the target during optics.
4) The optics plan is still under discussion. Min is working on snake and mudifi to verify the plan.
Min also show a few slices about that.
5) G2P work disk will be upgraded to 1T space from 14:30.
1) Work with Pengjia on HAPPEX DAQ project.
2) HAPPEX crate has been merged with the right HRS DAQ.
- Vince, Ryan, Melissa:
1) Took cosmic data for the left and right HRSs.
2) Left HRS, checked S1 PMT signals for after-pulsing issue. About four PMTs have a large amount of after-pulses, bars 5 and 6.
3) Found that the ADC amplitude of some of the PMTs is very low. Will compare to cosmic data in the Spring to find out if the low amplitudes existed then.
4) Initial conversation with Bogdan resulted in him requesting more plots for the ADC amplitudes. Melissa is making those plots now.
5) We probably need to replace the PMTs. However, Bogdan only has about half the PMTs for both S1 detectors (left and right).
(Those PMTs are out of production, we may need another solution.)
Check the details here.
1) Working on the coda scripts.
2) Updated the root to version 5.28 in the farm for analyzer. Ole has updated the root for adaq machines
to 5.30. Toby will update the root to this version too.
1) HAPPEX DAQ works for right HRS. Working on making another copy to left HRS. Will check the dead
time on the HRS DAQ due to adding the HAPPEX crate
9/21/2011: Present: Karl, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, Alex, Vince, Toby, JP, Yuxiang Zhao, Yi Zhou, Zhao Zhang
1) Copied HAPPEX DAQ (HARDWARE) into the right HRS and got it work.
2) Tungsten colorimeter is still under testing by Arne's student. We was told
that it need two more weeks to finish.
3) Show a few slices about the helicity signal that will be used during in g2p.
1) Working on 3rd arm DAQ. It is basically working right now. Doing the test with dE plane.
2) Worked with Whit(Engineering group) to design the support of the magnetic shielding.
The shielding material has been ordered.
1) Working on the simulation of C12, LHe, CH2 and LH2 with sieve slit in position. Trying to finish
the reconstruction code and verify if the we can separate the elastic peak of C12 from that of LHe.
Check here for details. This figure shows the CH2 elastic peak at 1.159 GeV beam energy.
The momentum elastic peak of helium and Carbon start to overlap after energy loss. Note that the momentum are not
reconstructed momentum. Smeared be the VDC resolution and reconstruction resolution, these 2 peak will overlap even more.
1) Working on mudifi to fit Geant4 sieve slit data for target field reconstruction routines
2) Tuning the snake field map
- Jie and Toby:
1) The target field measurement with translation table analysis note is ready for group review.
- Melissa, Vince and Ryan :
1) Working on HRS DAQ. The DAQs for both arm are basically working.
2) Melissa Updated start_coda script such that it starts left HRS DAQ on adaql1 and right HRS DAQ on adaql2
3) Took cosmic run with left HRS. Will take cosmic run for right HRS soon.
4) One rack keep losing power, Jack will take care of that
5) A couple HV channels were tripping for S2m, the HV card was suspected and replaced.
6) Ryan also looked into S1 and S2m TDC signal, some channels looks abnormal. Need to spend more time checking that.
1) Working on 3rd arm DAQ with Kalyan. Will do pedestal run soon. 2) Measure the sieve slits for hole positions. They are all match the design.
1)Working on test the intel CPU in DAQ.
9/14/2011: Present: Karl, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, Alex, Vince, Toby
1) Working on 3rd arm DAQ with Chao
2) The magnetic shielding (netic cylinder + co-netic foil) order has been placed
1) Shift schedule is ready to sign. Called for checking the description. Karl suggested not to
open the 2nd period for user now. For those who have signed shifts in the 2nd period we will honor
them the shifts they signed.
2) Working on the TOSP for septum field mapping
3) Compare SANE TOSCA map with our NEW TOSCA map together with Oscar. Could not find the reason.
4) Running simulation to check optics plan
1) Try to use intel CPU to replace the CPU board. 6~10 kHz DAQ rates is expected.
1) Working on HRS DAQ with Melissa and Ryan; Roc3 is now not running
2) Check the lumi HV crate, one channel keep tripping off. (Nothing we can do so far)
3) Plan to change the Left HRS HV operation interface as the one used for the right arm.
1) Prepare for optic plan
2) Working with Pendjia for helicity signal in the HAPPEX DAQ
1) Writing tech note for field measurement with translation table
2) Working with Pendjia for helicity signal in the HAPPEX DAQ. Made a comparison of time delay for 3 modeules.
1) Show the pre-shower|shower analysis result.
1) Working on left HRS ROC3. Took the structure fastbus interface from BigBite, but it does not work in roc3.
Will try to use an old CPU.
1) Updated the G2p wiki with README (instruction) for replay
2) Writing the tech note of the target field measurement on the can surface
1) Tested the dead time for various DAQ configurations. Check Elog entry 129 for details.
1) Working on 3rd arm DAQ with Kalyan
2) Glued the broken dE plane scintillator
9/7/2011: Present: Karl, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, JP, Alex, Vince,Toby, Yi Zhou, Yuxiang Zhao
1) The draft of run schedule is available here.
2) Discuss the safety and OSP|TOSP document issues
Welcome 2 USTC visiting scholar Yi Zhou and Yixiang Zhao. They will work on GEM and SOLID projects.
1) Agree to work on the TOSP for septum
2) Discuss the shift schedule and related things. Collaborators need to take 10 shifts to fulfill
the authorship requirement. Collaborators is not allowed to take TO shifts until got trained or
agree to receive the TO training well before shift starts.
3) Will talk to Oscar to find out the difference between old TOSCA and new TOSCA map for the target field. Here is the figure showing the difference between these models.
4) Design the magnetic shielding for the 3rd arm PMTs. Need to block 70 gauss field. Current chosen thickness, 0.062 inch, is over designed.
1) Reading tech notes and thesis. Prepare for optic plan.
1)Remind the student that helicity signal in the HRS is delayed
2)Setting up DAQ crate for the 3rd arm
3)Working on the magnetic shielding
1) Show his field analysis result
1) Not finalize his field analysis result yet, will show the result later
1) working on 3rd arm simulation
2) working on 3rd arm SC, made a holder and will glue it at the rest of this week
1) Ring buffer for HAPPEX DAQ works for helicity signal in the left arm. Will work on the right arm soon
8/31/2011: Present: Karl, Jixie, Jie, Melissa, Min, Chao, Pengjia, Kalyan, JP, Alex, Vince
1) Working on optics plan with Min. Need to recalculate e-C and e-H elas rates for 5.7 deg 2.5T.
2) Confirm septum polarity before Howard connect it to the power supply
3) Will perform B vs I curve for septum in middle Sep. Chao and Min will join, Toby too.
4) Karl suggest get (from Doug) the long cable which is used to connect to the hall probe.
5) Will simulate what thickness will be good for a mu-medal to shield low energy protons.
1) Working on the 3rd arm DAQ with Chao.
2) The dE plane of the 3rd arm is 6'(width) x 12'(height) in size, not 12 x 12. This is found
after it was opened. Chao will need to peel off the remainder glue and reglue it. Jixie will
check with Al about this change.
3) Kalyan agree to following replacing PMT for S1L2 (check minutes of Aug. 17 for detail).
1) Will join G2P and work on HRS DAQ. A 3rd crate might be added in order to improve DAQ rates.
2) Will following the issue of HV crate of Lumi. (check minutes of Aug. 17 for detail).
1) Working on optics plan with Jixie.
1) Working on the 3rd arm DAQ
2) open the SC with Kalyan. Will clean the opened SC and glue it back tomorrow
3) Need to run the simulation again for the 3rd arm using 6x24 size and latest design positions.
1) Checking on shower preshower. She will post her results to the elog.
1) Make a presentation about the DAQ plan for BPM|BCM.
After discussion we all agreed to put hardware into both left and right HRS crates such that each arm will have a complete DAQ system.
8/24/2011: Present: Karl, Jixie, Jie, Melissa, Ryan, Toby, Kalyan, JP, Chao, Min, Pengjia
1) The tosca model from SANE group was proved to be different from the one Robin recently
created. Check this plot to see how they vary. Both the field analyses from Jie and Toby also confirm that. Jixie will
talk to Robin to verify this.
2) Working on the Optics with and Min. A discussion with Jin Huang is ongoing.
3) The 3rd arm design is done. Karl mentioned that Al might have some small change in the design of
of the stand. Jixie will check this with Al.
4) Jixie will check the neutron background with the shielding with Poval.
1) Got some progrss on the HAPPEX DAQ:
A) The ring buffer works with helicity signal now, will move on to ADC signal
B) Still working on the TCPIP program
2) Compared the BPM|BCM signal of HAPPEX DAQ and that from the scaler. It looks like HAPPEX DAQ give
better resolution. Pengjia, Kalyan, Alex and Bob will make the final decision which one to choose.
3) Arne's student is testing the tungsten colorimeter with an external heat source in the Hall. Pengjia is watching on this. 4) There are 2 options to put BCM|BPM signal into the data stream:
option A: use socket program to insert the signal into the ET ring;
Option B: integrate the hardware into HRS crate;
Pengjia will compare these 2 options then make a choice.
1) Working on the optics. Join the discussion with Jin and Jixie. See her report here
1) Show the difference between old tosca map and new tosca map of the target field. Also show the
field analysis result. Check details here.
1) Showed his field analysis result using the new target field. The current ratio from his result
is consistent with Jie's result. But the position offsets still have some difference.
1) Writing a tech note about the design of the 3rd arm with Jixie
1) installed S1 back to HRS, will connect the cables and DAQ
1) Practice the calibration of shower and pre-shower using xgt2 data
1) Confirm that the trigger has been settled as the way we discussed in last meeting
2) Confirm that the 3rd arm DAQ can go up to 10KHZ
Present: Karl, Jixie, Vincent, Jie, Mellisa, Ryan, Toby, Doug, Alex, Kalyan, Zhihong
By Phone: Chao, Pengjia, Min
Discussion about the trigger
- Zhihong, Karl, Doug, Alex, Vince ... :
1) The HV of 3 channels in the (Left?) Cherencov are too high, need to be adjusted;
2) According to xgt2 data, the efficiency of S1 and S2 are still larger than 99%
3) The 2nd PMT in the Ledt S1 (S1_2_L) need to be replaced
4) S0 is already in the right HRS. We need to remove it
5) The HV of 2 channels of the Lumi crate (Slot 15, #7 and #8) should not be turned on,
otherwise they will cause the whole crate crashed;
After 3) and 4) fixed, the trigger for G2P will be the same as the regular HRS trigger.
- Chao, Min, Pengjia:
1) Reported from UVA that the target have been cooled down and with 5T field on Tuesday, they are trying
for 2.5T field today.
1) Bodgan told that S1m will not be ready for G2p;
2) Will organize rad.con.2 training and electrical lead training
3) As discussed with Jixie we need to place hall probe in the tunnel of both septa, encourage students
to help Jixie in measuring "B vs I" curve for septum coils;
1) Show the detection efficiency of the 3rd arm. It is about 85%.
Check here for details.
The following table is the rate of the 3rd arm for current design:
Please note that the detector efficiency is not considered yet.
2) Both of the 2 target field analysis from Jie and Toby shows that the current ratio is closed to 1.15. Jixie will
talk to survey group for checking the position for mistakes. Jixie also suggested Jie and Toby to compare the new target
field map from Robin to the exist one. If we can not explain this 1.15 current ratio, we need to measure the field on the can surface again.
3)Will measure the "B vs I " curve for septa when power supply ready. We need 2 gauss meters and 2 probes to place in the septa
1) Found that the current ratio is about 1.15. Will check bug for Toby and try to improve the field analysis
1) Found that the current ratio is about 1.14.
2) Will check bug for Jie's program and try to improve the result
1) Checking the ADC of preshower and shower
1) Will look into the HV problem mentioned above.
2) Check the Y distribution at S1m to see if S1m will lose acceptance. This position is projected
from the focus plane. See the details here.
Note: we will not be able to use S1m.
3) Will clean the unnecessary cables for the HRS ;
Present: Karl, Jixie, Chao, Pengjia, Min, Vincent, Jie
By Phone: Mellisa, Ryan, Toby
1) Target is not cool down yet, still filling helium. Ryan, Toby and Mellisa are already there.
2) Can not find 2 cm SC
3) Got approved that we will have 12 days of 1.7 GeV beam (may be a old news to someone).
4) Tried to extend to run the experiment during Christmas
1) Talked to Bogdan and Albert for 2cm SC. There is no EXTRA 2 cm SC that can be used for us.
2) Talked to Brad and Bob about 10K Hz DAQ rate with ADC. Both of them are said there need some
efforts to make this happened.
3) Finalized the 3rd arm design:
A) Since 2cm SC is not available, we will switch back to 1 cm SC for the 1st plane;
B) The neutron shielding will be 30 cm thick and we will use the exist borated plastic;
C) The width of of SC plane will be extended from 6 to 8 inches;
D) We will put a removable 2mm aluminum shielding to shield low energy background from the target.
We will not use this shielding for 1.2 GeV. For 1.7 and 2.3 GeV, we will use it only if the DAQ rate
can not reach 10k Hz
E) We changed the detected energy range for 1.7 and 2.3 GeV.
4) Calculated the maximum raster size (requested by Karl). It is ok to extended to 11 mm.
raster size assuming 25.4mm target cell diameter.
We will have the target cell diameter of 27mm, , which will allow to
extend to 11.5 mm.
1) Simulated the 3rd arm. There is a bug in calculating the asymmetry. Need to recalculate the time
2)After fixed the bug in the code, found the final position and rates:
Check details here.
1) Solve the snake problem in linux by using -fno-automatic option.
2) Wrote a note about the optics discussion with Jixie.
1) Will looking to the low current data taken during Xgt2 to check noise of adc in BCM|BPM
2) Arne will test the tungsten colorimeter with heater some time next week
3) The ring buffer is working, will move on to read ADC and helicity signal, and TCPIP program which
will be used to communicated between the ring buffer and HRS DAQ
1) Shown improved result. After optimizing the fitting ranges, the variation of data to tosca reduce to 3%,
and the position offsets are all within 1 cm. Check | this for details.
8/3/2011: Present: Karl, Jixie, Chao, Pengjia, Mellisa, Vincent, Ryan, Toby, Jie
1) Organized student go to UVA to participant the target cool down
1) The detector response function is ready in the Geant4 simulation program.
| Energies deposited by elas proton
| Hits in neutron simulation
| Total energy deposited by neutron. (Please note that the top panel is still for the 2nd plane, not the 1st plane.)
The energy deposited into the 2nd plane indirectly from the neutron is about 188 MeV per 100K 2.3 GeV electrons
or 160 ns with 100 nA beam current. It turns out we must have the neutron shielding!
2) To shield the neutron, 20 cm thick plastic (Polyethylene) can block 80~85% of the neutron flux.
Check detail here: 
3) Third arm detector design: Al confirm that we can put 20 cm thick plastic shielding.
The detector will be located 70 degrees, 2.1 meter away from the target. The SC plane size is 6'X 24',
with rotation by the vector (r=1, theta=70, phi=0) about 36 degrees (for 5T target field) or 18
degrees(for 2.5T target field). The 60 degrees of rotation about Y axis is not necessary since the shielding
for neutron is in place. The first layer of SC will be 2cm, and the 2nd layer will be 5 cm or 2 inches.
We need to find 2cm scintillator. The center position for each beam energies are the following:
1.159 Horizontal_offset = 6 cm, Vertical offset = 0 cm OR Horizontal_offset = 0 cm, Vertical offset = 18 cm
1.706 Horizontal_offset = 0 cm, Vertical offset = -21 cm
2.257 Horizontal_offset = 0 cm, Vertical offset = -15 cm
The position can allow 1~2 cm off, and the angle can allow 2 degrees off.
4) Have already inform Al with this design.
1) Ran Geant4 simulation for the 3rd arm. Verified the rotation and position.
2) Recalculate the rates. It will take about 14, 10, 7 hours to measure Pb*Pt to 5% uncertainty with 1.2,
1.7 and 2.3 GeV beam. If we increase the SC wider it will shorten the time for 1.2 GeV but longer the time for
the others due the the signal to background noise get worse. We still need to optimize the design.
1) Working on trigger on S1m
2) Looked in the Y distribution of S0, it is about 15 cm, it seems ok for acceptance. | Figure 1 | Figure 2
3) Got the timing module working after replacing the broken cables. Bob suggested to unplug HRS cable and reconnect them.
4) Will work on the HOWTO wiki pages
1) Checked the probe with known field in the magnetic group with Jixie, Chao and Jie. Both the 3-axis probe and
the axial probe seems within 1% accuracy. But we did not check the variance of the axial probe in rotation.
Plan to do this in UVA.
2) Determined the current ratio to be around 1.0. Jixie suggested to check again.
1) Improved the field analysis result. He analysis the data X_table=44 or X_table=46, which is most closed to the z axis.
The new current ratio is 1.151, the position offsets are Xoff~=1.2 cm, Yoff=-0.5 cm, Zoff=0.4cm. Rotation and Xoff is not
sensitive at all in this data set because the field is almost flat when x is closed to 0.
Check the detail here: http://hallaweb.jlab.org/experiment/g2p/collaborators/jie/fieldmap.pdf
2) Will use large X data to fit for rotation and Xoff.
1) Tried to pull fortran routines from cern packages. But too many to pull! Jixie suggested stop doing this.
1) Successfully build the ring buffer to store BCM|BPM signal on HAPPEX DAQ
2) Will post the target connection diagram which can help to understand the target cool down. Check Elog entry 124
for detail https://hallaweb.jlab.org/dvcslog/g2p/124
7/27/2011: Present: Karl, Jixie, Chao, Pengjia, Mellisa, Vincent, Ryan, Toby, Jie
1)Neutron flux simulation.| details
There will be about 1.3~1.5 neutrons for each 2.3 GeV electron hitting the local dump.
For 6'x24' 3rd arm detector located at 70 degrees and 2.1 meters away from the target,
there will be about 0.0012 neutrons per 2.3 GeV electron. If converted to rates,
it will be about 800 MHz @ 100 nA beam current. This result has been confirm by a FLUKA program ran
by Lorenzo Zana (who is checking the radiation for PREX2).
Jixie will work on shielding the neutron. 2)Working on third arm simulation. |details
If rotated the 3rd arm by Y axis with 60 degrees (the the SC plane will be 10 degrees with
respected to X axis), we can have 3 advantages:
A)reduce the neutron flux by a factor of 5 (=sin(70deg)/sin(10deg))
B)do not need to cut the scintillator plane
C)effective thickness doubled and will have better ability to cut ADC for triggers. (need to simulate ...)
But it require more space in the target platform. Need to check with Al
3)Can not understand how the neutron will affect the 3rd arm trigger without the detector response
simulation. Jixie will add detector response into the Geant4 simulation.
1) Working with Jixie on the 3rd arm simulation. Confirm the rotation angles and tried to find
the best positions for each beam energies.
2) Recalculated the rate for the 3rd arm.
1) Tried to improve the analysis on the target field
1) Checked the self consistency of the HRS transportation packages.
2) The SNAKE works only in SunOS but not. MUDIFI works on jlabl2 but not on jlabs3. It was assumed to
be the cernlib problem. Karl suggested to pull the needed fortran routines out of cern packages and
build a stand-alone snake program.
1) Finished checking the Cherencov.
2) Will move on to check the ADC signal of pre-shower and shower.
1) Checking timing for L1A (Level 1 acceptance) 2) Working on the trigger on S1m
1) Optimized the fitting range in the field analysis
7/20/2011: Present:Karl, Jixie, Chao, Pengjia, Mellisa, Vincent, Ryan, Toby, Alexandre, Jie Liu
1) The design of the 3rd arm:
A) We need to place the 3rd arm at 70 degrees, not 72. We also have to move the detector
from 1.5 meter to 2.1 meter such that the local dump itself could block some background.
B) | Figure 1 shows the hit positions on the SC plane. The distribution look like a rotated strip.
In order to select a small range of theta, the SC need to be rotated to some angle.
C) | Figure 2 and | Figure 3 show the hit position for protons within and out of the theta range.
Please note that these protons are generated in a flat distribution of theta and Energy.
they are not elastic proton. Jixie will run with elastic proton again to check the rotation.
D) We might need to cut the SC from 12 inches width to 6 inches.
E) JP mentioned that it might help if the SC plane be rotated about Y axis by 70 degrees.
With this rotation the whole SC will have minimum cross section area facing the local dump.
2) Still working on neutron flux simulation.
1) Simulate the 3rd arm detector located at various distance from the target (1.5, 1.75,2.0, 2.25 meter).
The detector size in the simulation is 6' x 24'. According to his result the dilution factor is about 17%
and the total time to accumulate 5% accuracy data is about 9 hours.
1) Check the no target field HRS transportation code
1) learning the HAPPEX DAQ
1) Show field analysis result on the can surface data set. After improve, the variation of measured field
to the map are within 1%. But the position offsets are still in 2 cm level. This analysis used only point A
position to calibrate all the data set, which leaves room for further improve. Jixie ask Toby to check
the code carefully.
1) Show field map analysis. His result does not agree with Toby's.
1) Working on mapping the HRS HV. Finish ROC3 and ROC4. Check Elog entry 122 for detail:
2) Will work on DAQ with Alex
3) Will check PMT signal for S1 and S2m
1) Working of the safety documents which is drawn from SANE group.
2) Checking the Cherencov, found the ADC for 1 photon-electron, but this result still need to be improved.
7/13/2011: Present:Jian-ping, Jixie, Chao, Pengjia, Mellisa, Vincent, Ryan, Toby, Alexandre, Jie Liu, Karl
1) Continue checking the Cherencov detector. Applied geometry cut to the ADC. She need to find out
the ADC value that 1 PHE corresponding to. Check her plots in the Elog, Item 120:
1) Showed the preliminary result for the field map analysis with the can surface data set. Most data
points are within 6% variation. But the position offsets are as large as 2 cm.
1) Showed his result of the field map analysis. Most data points are with 10% variation to the
TOSCA map. Jie's result does not agree with Toby's in position and current ratio. Both of them are
requested to check their results more carefully.
1) Tried to build ring buffer in HAPPEX DAQ. Will do the tungsten calorimeter calibration in August.
1)Looked into the HV for HRS
2)Finish the splitter box. Check Elog entry 121 for detail: https://hallaweb.jlab.org/dvcslog/g2p/121
1) Tried to learn and run the Geant4 simulation for the 3rd arm.
1) Looking into the neutron background data provided by Pavel. 2) Designing the 3rd arm.
7/6/2011: Present: Jian-ping, Kalyan, Jixie, Chao, Pengjia, Mellisa, Vincent, Ryan, Toby, Alexandre, Jie Liu
1) Discussed the run plan with QWEAK group, ask for 12 days for 1.7 GeV. Not be granted yet.
Currently Hall C allow only 5 days, or If they are down then we can take the time for 1.7 GeV.
The plan is still under discussion.
2) If Hall B is not available, we need to be able to run 1.2 GeV. We should be very flexible and ready for changes.
3) Karl has the readiness review available. Everyone is encouraged to read and comment.
4) JP will leave from 7/22 to 8/17. Kaylan will leave to India from 7/8 to 7/29.
Alex will on vacation from 7/23 to 8/11. Karl will be on site from 7/11.
1) Just back from the summer school. Will continue working on the HRS package issues.
1) Taking the detector geometry into account, recalculated the background of the 3rd arm using the epc code.
2) Discuss the 3rd arm design: 12 inch width will accept +/-4 degrees in theta. With such a wide
range it will take about 16 hours to measure Pb*Pt to 5% accuracy. It is too long ...
3) Finish the target field measurement with Jixie and Toby.
1)Leaning HAPPEX DAQ, will use it for BCM and BPM
2)will do calorimeter test. The tungsten calorimeter will eventually used to calibrate BCM.
3)Learning how the Harp work and how to use it to calibrating BPM.
1) Finish Target filed mapping with Jixie and Chao.
2) Working on analyzing the field mapping data.
1) Working on Left HRS trigger
2) Will check the material of the target with Josh
1) Checking the Cherencov detector in HRS. Tried to find how many photon-electron (PHE) each PMT has
in the Xgt2 data set. She need to apply the geometry cut and try again.
1) Analyzing the filed measurement data.
1) Just Finished the field measurement with help from Chao and Toby. Pengjia and Jie also provide appreciable help.
2) Measured the field in 2 ways. One is measuring the field on the can surface with an axial probe (borrowed
from Josh) fixed in a aluminum block.
Another is measuring the filed with a 3-axis probe on the translation table.
The survey group came to the target lab and survey some positions. Just got the survey data yesterday.
Jie and Toby are working on analyzing the field map measurement. Jie is focus on translation table data set
and Toby is working of the can surface data.
3) Talk to Pavel in the Rad. Con. group for the neutron background simulation. Pavel will provide
the neutron distribution and model information such that Jixie will be able to work on radiation calculation.
6/22/2011: Present: Kalyan, Jixie, Chao, Pengjia, Mellisa, Vincent, Ryan, Toby, Alexandre, Jie Liu
1) Tested 4 cables for the slow raster with Ryan
2) Worked with Chao and Jixie to prepare to measure the target field
1) Continue working on the HRS trigger
1) Replaying xgt2 data. Looking at the pedestal of adc to find out if there are dead channels
2) Updating the safety document for g2p based on the one used in SANE experiment
1) Preparing a computer in the hall for the DAQ of the 3rd arm 2) Preparing for the target field measurement
1) Continue working on the remote control program for the function generator
1) Continue designing the 3rd arm
2) Got confirmed from Al that 3rd arm can not be used in GEP
1) Learning the optics from Nilanga, will meet with Vince to discuss the optics plan
2) Tried to test the magnetize effect of the translation table. Replaced some steel bolts and nuts
with plastic ones. Toby and Chao will make a coil to test the stainless steel.
3) Running simulation to get local beam dump opening for Al. But the HRS transportation routines from Min do not work.